2 times one number added to 3 times another number is 54. The second number is 2 less than the first number.
Answer by darq(90)

#1 number = x
#2 number = y

2x+3y=54
x=y+2

2x=54-3y
2x=2y+4

2y+4=54-3y
5y=50
y=10
so x is 2 greater than y which means x is 12 :)